Compound X has a molar mass of 135.12·gmol−1 and the following composition:

element/mass

carbon: 44.44%

hydrogen: 3.73%

oxygen: 51.83%

Write the molecular formula of X

1
Expert's answer
2021-10-24T01:24:57-0400

v(C)=135/44.44=3mol

v(H)=135/3.73=36mol

v(O)=135/51.83=3mol

C3H36O3